Minutes of the Linac4 Diagnostics Working Group Meeting held on 3 November 2008

Present: D.Gerard, G.Tranquille, L.Soby, C.Dutriat, E.Bravin, B.Mikulec.

Agenda:

1. Communications

2. Follow-up of open actions

3. Assembly status of the emittance meter (D.Gerard)

4. AOB

1. Communications

K.Hanke transmitted the information that the IPHI diagnostics line will not be delivered to CERN. Therefore this will become a first priority item for the coming months. U.Raich should be able to come up with more information at the next WG meeting as there has been a crisis meeting with the French last week where he was present.

2. Follow-up of open actions

-. Emittance meter: E.Bravin confirms that the -1 kV power supply for the FC and the +300 V power supply for the SEMgrids is available. It is supposed that the ADC for the FC will be the one developed by CERN/BI and currently under test with controls software being prepared by A.Lokhovitskiy. C.Dutriat volunteered to take follow up all cabling issues connected with the emittance scanner.

3. Assembly status of the emittance meter

D.Gerard presented the assembly status of the emittance meter (see montageE-meter.pdf). A 3D model exists and will be transferred to the TS archive. In terms of hardware the potentiometer for the movement of the SEMgrid has been installed (0.75 mm wire spacing). All the hardware pieces except the cabling are prepared for assembly in the lab in building 865 where also measurements with LEDs will be performed.
This week the integration outside the vacuum tank is planned and for the week after the assembly of all the pieces into the tank. C.Dutriat will help to provide the cabling and infrastructure to perform the alignment.

It will be possible to adjust the angular position of slit and SEMgrid as foreseen. The program for the PLC control (movement control) is ready except some details that still have to be sorted out.

For higher beam energies it will be necessary to add a cooling system for the slit. Slit cooling, slit dimensions, the separation slit - SEMgrid and possible optics in-between are subject to simulation studies that will be performed by the new fellow. In any case the slit will have to be exchanged as well as a new tank designed and produced.

R.Scrivens confirmed after the meeting that the final installation of the emittance meter will not be required before February/March 2009. Source measurements with the emittance meter will only be performed after the measurements with the Faraday cup alone.

-- BettinaMikulec - 03 Nov 2008

Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf montageE-meter.pdf r1 manage 935.7 K 2008-11-03 - 18:01 BettinaMikulec Assembly status of the emittance meter (D.Gerard).
Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2008-11-04 - BettinaMikulec
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright